a ballon move with constant acceleration 6m/s2ย ย in upward direction after 3 sec a particle drop from balloon .find thhe distance between balloon and the earth when particle reachย  on the groundย 
ย 

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

At 3 sec height and speed of balloon
h = 27 m
v = 18 m/s upward

Height of partical after being dropped = 27 + 16.2 = 43.2 m

Time req = t up + t down = 1.8 sec + 2.9 sec

Distance traveled by balloon = 0.5 *6 * 6.7*6.7

= 134.67 m

Regards